Simultaneous Equations: sorry didnt see Cat. for that:
Solve for x,y, and z:
x + 2y + z = 2
5x = z
y = -2x

Answer by mananth(16946)

x + 2y + z = 2
5x = z
y = -2x
substitute the value of z & y in first equation
x+2*(-2x)+5x=2
x-4x+5x=2
2x=2
/2
x=1
..
y=-2
..
z=5